OpenStudy (anonymous):

4x=16? The x is raised?

OpenStudy (anonymous):

x+2

OpenStudy (anonymous):

im sorry x =2

OpenStudy (anonymous):

What? x = 4

OpenStudy (saifoo.khan):

x = 16/4 x = 4

OpenStudy (anonymous):

the x is raised

OpenStudy (anonymous):

it's not multiplied

OpenStudy (anonymous):

If it's raised, take logs of both sides

OpenStudy (anonymous):

4 squared is 16 so x=2

OpenStudy (anonymous):

Carolyn, what kind of math are you doing?

OpenStudy (anonymous):

even if you log both sides you still get 2

OpenStudy (anonymous):

If the x is raised, yeah, it equals 2...okay

OpenStudy (anonymous):

The original poster should provide proper symbology when posting problems

OpenStudy (anonymous):

the original poster did..

OpenStudy (anonymous):

proper symbology would be 4^x = 16

OpenStudy (anonymous):

4x=16? The x is raised it doesn't really matter we get the point don't we?

OpenStudy (anonymous):

\[4^x=16\] \[x=2\] since \[4^2=16\]

OpenStudy (anonymous):

take the log???!!! no thanks

OpenStudy (anonymous):

bravo